### v2

# A Representation Transformation of Parametric Feynman Integrals

Abstract
A transformation on homogeneous polynomials is proposed, which is further applied to parametric Feynman integrals. The two representations related through this transformation are dual to each other. And it naturally leads to dualities of Landau equations and linear integral relations between the two representations. For integrals with momentum-space correspondences, the dual representation is equivalent to the Baikov representation.
I Introduction
Feynman integrals are a crucial element of perturbative quantum field theory. There are several representations of Feynman integrals, with the most well-known being the Feynman-parameter representation [1, 2, 3, 4]. Variants of the Feynman-parameter representation include the Lee-Pomeransky representation [5, 6] and the version used by the author in refs. [7, 8, 9, 10]. The parametric representation is powerful in the sense that it allows for both direct integration [11, 12, 13, 14, 15] and integral reduction [16, 6, 17, 7, 8, 9, 18, 19, 20], and Landau equations [21, 22, 23] can easily be formulated within this framework (see e.g. ref. [24]). Furthermore, the parametric representation can reveal some hidden mathematical structures of Feynman integrals due to its connections with hypergeometric functions [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], periods [36, 37, 14], coaction [38, 39, 40, 41, 42], etc.
Another widely used representation is the Baikov representation [43, 44, 45, 46, 47] and its variants [48, 49, 50]. Although it is less suited for direct integration, the Baikov representation has the advantage that implementing cuts is straightforward [48, 51, 52], and it is more compatible with twisted cohomology theory [53, 54, 55, 56]. The Baikov representation has been extensively applied to integral reduction [57, 58, 53, 54, 55, 56, 59, 60], constructing differential equation basis [61, 62, 63], Landau analysis [64, 65, 66, 67], etc..
While it is evident that these two representations should be equivalent, since both are derived from the momentum representation, it was not clear how to directly derive one representation from the other. This problem is addressed in this article. It is demonstrated that these two representations are related by a simple transformation. A transformation on homogeneous polynomials is introduced in section II. The transformation is then applied to parametric Feynman integrals in section III. It is proven that the two representations related through this transformation are dual to each other. The equivalence between the dual representation and the Baikov representation is proven in section IV.
II Polynomial transformation
We consider the following transformation on a homogeneous polynomial $\mathcal{F}(x_{1},x_{2},\dots,x_{n+1})$.
$u_{i}\equiv\frac{1}{\mathcal{F}}\frac{\partial\mathcal{F}}{\partial x_{i}},$ (II.1a)
$\mathcal{G}(u)\equiv\left.\frac{1}{\mathcal{F}}\right|_{x=x(u)}.$ (II.1b)
We assume that the degree of $\mathcal{F}$ is $L+1$. Then obviously
$\sum_{i=1}^{n+1}u_{i}x_{i}=L+1.$ (II.2)
Taking the derivative of both sides with respect to $u_{i}$, we get
$x_{i}+\sum_{j=1}^{n+1}u_{j}\frac{\partial x_{j}}{\partial u_{i}}=0.$
On the other hand,
$\begin{split}\frac{1}{\mathcal{G}}\frac{\partial\mathcal{G}}{\partial u_{i}}=\mathcal{F}\frac{\partial}{\partial u_{i}}\left(\frac{1}{\mathcal{F}}\right)=-\frac{1}{\mathcal{F}}\sum_{j=1}^{n+1}\frac{\partial\mathcal{F}}{\partial x_{j}}\frac{\partial x_{j}}{\partial u_{i}}=-\sum_{j=1}^{n+1}u_{j}\frac{\partial x_{j}}{\partial u_{i}}.\end{split}$
A combination of the above two equations leads to
$x_{i}=\frac{1}{\mathcal{G}}\frac{\partial\mathcal{G}}{\partial u_{i}}.$ (II.3)
Thus we see that the transformation defined by eq. II.1 is reciprocal.
Generally speaking, the transformation defined by eq. II.1a may not be reversible. Even if it is, the obtained function $\mathcal{G}(u)$ may not be a rational function. However, for the $\mathcal{F}$ polynomial of a loop integral with a complete set of propagators, $\mathcal{G}$ is indeed a rational function, as will be clear in the subsequent sections.
III Representation transformation
As was shown in refs. [7, 8, 9], a Feynman integral can be parametrized by scalar integrals of the following form.
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})\\ =&\frac{\Gamma(-\lambda_{0})}{\prod_{i=m+1}^{n+1}\Gamma(\lambda_{i}+1)}\int\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\prod_{i=1}^{m}x_{i}^{-\lambda_{i}-1}\prod_{i=m+1}^{n+1}x_{i}^{\lambda_{i}}~{}.\end{split}$ (III.1)
[…]
We introduce the function [9]
$w_{\lambda}(x)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}v~{}\frac{1}{(v-i0^{+})^{\lambda+1}}e^{ivx}~{}.$ (III.2)
We have
$w_{-1}(x)=$ $2\pi\delta(x)~{},$
$w_{\lambda}(x)=$ $2\pi\theta(x)\frac{x^{\lambda}}{\Gamma(\lambda+1)},~{}\lambda\geq 0~{}.$
And we have
$\frac{1}{x^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}v}{2\pi}~{}w_{\lambda}(x)e^{ivx}~{}.$ (III.3)
Obviously we can write $\frac{1}{\Gamma(\lambda+1)}\int_{0}^{\infty}\mathrm{d}xx^{\lambda}$ as $\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}xw_{\lambda}(x)$. By expressing $w_{\lambda}(x)$ (or $x^{-\lambda-1}$) in terms of $v^{-\lambda-1}$ [or $w_{\lambda}(v)$], it is easy to get
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}v_{i}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\int_{-\infty}^{\infty}\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\int_{-\infty}^{\infty}\mathrm{d}\Pi_{v}^{(n+1)}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},\end{split}$ (III.4)
where the last step can be proven by inserting an integral $\int\mathrm{d}s\delta(s-\mathcal{E}^{(1)}(v))$, rescaling $x$ and $v$ by $x_{i}\to\frac{x_{i}}{s}$ and $v_{i}\to sv_{i}$, and finally integrating out $s$.
We consider a special class of integrals of which the $\mathcal{F}$ polynomials are complete. Here, by complete, we mean that there are $n$ linearly independent matrices $\Lambda_{i,jk}$ such that
$\sum_{j,k=1}^{n+1}\Lambda_{i,jk}u_{j}x_{k}=0,~{}i=1,2,\dots,n~{}.$ (III.5)
We further assume that
$\Lambda_{i,j(n+1)}=$ $0,~{}j\neq n+1~{},$ (III.6a)
$\sum_{j=1}^{n}\Lambda_{i,jj}=$ $n_{0}\Lambda_{i,(n+1)(n+1)}~{},$ (III.6b)
with $n_{0}$ a constant. As can be seen in appendix B, normal loop integrals do satisfy these conditions. The first condition is a consequence of the fact that $\mathcal{F}$ is linear in $x_{n+1}$. For a complete homogeneous $\mathcal{F}$ polynomial satisfying the conditions in eq. III.6, it can be proven that (see appendix A)
$\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)=C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}^{\prime}}~{},$ (III.7)
where $\lambda_{0}^{\prime}\equiv\lambda_{0}+\frac{n-n_{0}}{L+1}$ and $C_{\mathcal{F}}$ is a constant. Thus we get
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)C_{\mathcal{F}}\\ &\times\int_{-\infty}^{\infty}\mathrm{d}\Pi_{u}^{(n+1)}\mathcal{G}(u)^{-\lambda_{0}^{\prime}}\prod_{i=1}^{m}w_{\lambda_{i}^{\prime}}(u_{i})\prod_{i=m+1}^{n+1}u_{i}^{-\lambda_{i}^{\prime}-1}\\ \equiv&(2i\pi)^{-n-1}\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)C_{\mathcal{F}}\frac{\Gamma(-\lambda_{0})}{\Gamma(\lambda_{0}^{\prime})}\\ &\times I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})~{},\end{split}$ (III.8)
where $\lambda_{n+1}^{\prime}=\lambda_{n+1}+n_{0}+1$, and $\lambda_{i}^{\prime}=\lambda_{i},~{}i\neq 0,n+1$.
Thus, we have proven that $I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})$ equals $I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ up to some constants, and there is an elegant duality between these two representations, which is summarized in table 1.
$\mathcal{G}=\frac{1}{\mathcal{F}}$ $\mathcal{F}=\frac{1}{\mathcal{G}}$
$u=\frac{1}{\mathcal{F}}\frac{\partial\mathcal{F}}{\partial x}$ $x=\frac{1}{\mathcal{G}}\frac{\partial\mathcal{G}}{\partial u}$
$\frac{1}{u^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{2\pi}~{}w_{\lambda}(x)e^{iux}$  ${}w_{\lambda}(x)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}u~{}\frac{1}{u^{\lambda+1}}e^{iux}$
Table 1: The duality between the two representations.
IV The Baikov representation
We consider a loop integral with a complete set of propagators
$v_{i}=\sum_{j,k}A_{i,jk}l_{j}\cdot l_{k}+2\sum_{j,k}B_{i,jk}l_{j}\cdot p_{k}+C_{i},\quad i=1,~{}2,\dots,n~{},$ (IV.1)
where $l$ are the loop momenta and $p$ are the external momenta. For simplicity, I assume that all the propagators are linearly independent and the Gram determinant of the external momenta $p$ does not vanish. For future convenience, we collectively denote $l$ and $p$ by $q$. Scalar integrals can be parametrized by integrals of the form [44, 45]
$J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})=\int\prod_{i=1}^{n}\mathrm{d}v_{n}~{}P^{-\lambda_{0}-\frac{1}{2}(L+E+1)}\prod_{i=1}^{n}v_{i}^{-\lambda_{i}-1}~{},$ (IV.2)
where $E$ is the number of the external momenta $p$. Here we have omitted some irrelevant prefactors for brevity. The polynomial $P(v)$ is proportional to the Gram determinant of the momenta $q$. That is
$P(v)\equiv\frac{\det(q_{i}\cdot q_{j})}{\det(p_{i}\cdot p_{j})}~{}.$ (IV.3)
In this section, we will prove that $J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})$ is equivalent to the integral $I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ in eq. III.8 up to some constants.
Taking $\mathcal{E}^{(1)}=|u_{n+1}|$ and $m=0$ in the integral $I_{u}$, we get
$I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$
$=$ $\Gamma(\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\left[\mathcal{G}_{+}^{-\lambda_{0}^{\prime}}+\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n}^{\prime}+1)}\right]\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
$=$ $\Gamma(\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
×$\times\left[\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{-i\pi(\lambda_{n}^{\prime}+1)}+\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n}^{\prime}+1)}\right]$
$=$ $2\Gamma(\lambda_{0}^{\prime})\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
$=$ $\frac{-4i}{\Gamma(1-\lambda_{0}^{\prime})}\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)$
×$\times\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\theta\left(-\mathcal{G}_{-}\right)\left(-\mathcal{G}_{-}\right)^{-\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}~{},$
where $\mathcal{G}_{+}\equiv\left.\mathcal{G}\right|_{u_{n+1}=1}$ and $\mathcal{G}_{-}\equiv\left.\mathcal{G}\right|_{u_{n+1}=-1}$. Due to the positivity of $|u_{n+1}|$, the integral recieves contributions from both $u_{n+1}=1$ and $u_{n+1}=-1$. The second step is carried out by reversing the path of integration for all $x_{i}$ for the first term in the squared bracket. And in the last step, we have closed the contour around the branch cut. Notice that we use such a convention that $\mathrm{Im}\{\mathcal{F}\}<0$. Thus $\mathrm{Im}\{\mathcal{G}\}>0$ when $\mathrm{Re}\{\mathcal{G}\}<0$. Now the region of integration is constrained by the theta function $\theta\left(-\mathcal{G}_{-}\right)$. Since $n_{0}=n-\frac{1}{2}(L+1)(L+E+1)$ (see appendix B), we have $\lambda_{0}^{\prime}=\lambda_{0}+\frac{1}{2}(L+E+1)$. Comparing the above integral with the Baikov integral in eq. IV.2, we see that these two integrals are equivalent if
$P(v_{1},v_{2},\dots,v_{n})=-\mathcal{G}_{-}=-\mathcal{G}(v_{1},v_{2},\dots,v_{n},-1)~{}.$ (IV.4)
Immediately, we will see that this is indeed true.
We define
$A_{ij}\equiv$ $\sum_{k=1}^{n}A_{k,ij}x_{k}~{},$
$B_{ij}\equiv$ $\sum_{k=1}^{n}B_{k,ij}x_{k}~{},$
$C\equiv$ $\sum_{i=1}^{n}C_{i}x_{i}~{}.$
Then we have [7]
$\mathcal{F}=\sum_{i,j,k,l}p_{k}\cdot p_{l}\tilde{A}_{ij}B_{ik}B_{jl}+(x_{n+1}-C)U~{},$ (IV.5)
where
$U\equiv$ $\det(A_{ij})~{},$
$\tilde{A}_{ij}\equiv$ $\left(A^{-1}\right)_{ij}U~{}.$
Let $a_{ij}$, $b_{ij}$, $\alpha_{ij,k}$, and $\beta_{ij}$ be those defined in ref. [8]. That is, they are solutions to the following equations.
$\sum_{l}a_{il}B_{l,jk}=0,$ (IV.6)
$\sum_{l}b_{il}A_{l,jk}=0,$ (IV.7)
$\sum_{kl}\alpha_{ij,k}a_{kl}A_{l,mn}=\frac{1}{2}\left(\delta^{im}\delta^{jn}+\delta^{in}\delta^{jm}\right),$ (IV.8)
$\sum_{kl}\beta_{ij,k}b_{kl}B_{l,mn}=\delta^{im}\delta^{jn}~{}.$ (IV.9)
Then it can be proven that (see appendix B)
$\begin{split}\mathcal{G}=&u_{n+1}\det\left[\sum_{kl}\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\right.\\ &\left.+u_{n+1}\sum_{mn}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{},\end{split}$ (IV.10)
where $u_{i}$ are those appeared in eq. II.1, $g_{ij}$ is the inverse of the Gram matrix $p_{i}\cdot p_{j}$, and
$B^{\prime}_{ij}(u)\equiv\frac{1}{2u_{n+1}}\sum_{k,l}\beta_{ij,k}b_{kl}\left(u_{l}+C_{l}u_{n+1}\right)~{}.$
On the other hand, $q_{i}\cdot q_{j}$ can be understood as a block matrix. Then we have
$P=$ $\frac{1}{\det(p_{i}\cdot p_{j})}\det\begin{pmatrix}l_{i}\cdot l_{j}&l_{i}\cdot p_{j}\\ l_{i}\cdot p_{j}&p_{i}\cdot p_{j}\end{pmatrix}$
$=$ $\det\left(l_{i}\cdot l_{j}-\sum_{mn}g_{mn}l_{i}\cdot p_{m}l_{j}\cdot p_{n}\right)~{}.$
It is easy to get
$l_{i}\cdot p_{j}=$ $\frac{1}{2}\sum_{k,l}\beta_{ij,k}b_{kl}(v_{l}-C_{l})\equiv-B^{\prime\prime}_{ij}=-\left.B^{\prime}_{ij}(v)\right|_{v_{n+1}=-1}~{},$
$l_{i}\cdot l_{j}=$ $\sum_{k,l}\alpha_{ij,k}a_{kl}(v_{l}-C_{l})~{}.$
Thus
$P=\det\left(\sum_{kl}\alpha_{ij,k}a_{kl}(v_{l}-C_{l})-\sum_{mn}g_{mn}B^{\prime\prime}_{im}B^{\prime\prime}_{jn}\right)~{}.$ (IV.11)
Comparing eq. IV.10 with eq. IV.11, we get eq. IV.4.
V Discussion
A reversible transformation on homogeneous polynomials is proposed. The transformation is further applied to parametric Feynman integrals, since a parametric integral is characterized by a homogeneous polynomial. An exact duality between the two representations related through this transformation is proven, as summarized in table 1. Furthermore, it is proven that the dual representation is equivalent to the Baikov representation.
The transformation and the corresponding duality can be applied to various aspects of Feynman integrals, including Laundau analysis and integral reduction. Specifically, the transformation in II.1 establishes a duality between the first (second) Landau equation in the parametric representation and the second (first) Landau equation in the Baikov representation. Additionally, eq. III.5 establishes a correspondence between a syzygy equation in the Baikov representation [57, 58] and a parametric annihilator in the parametric representation [6, 17].
Acknowledgements.
This work is supported by the Natural Science Foundation of China (NSFC) under the contract No. 11975200, and by Guangdong Major Project of Basic and Applied Basic Research (No. 2020B0301030008).
Appendix A Proof of eq. (III.7)
Combining the homogeneity condition eq. II.2 with those conditions in eqs. III.5 and III.6, we have
$0=$ $\sum_{j,k=1}^{n+1}\Lambda_{i,jk}u_{j}x_{k}$
$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda_{i,jk}u_{j}x_{k}+\Lambda_{i,(n+1)(n+1)}u_{n+1}x_{n+1}$
$=$ $(L+1)\Lambda_{i,(n+1)(n+1)}$
$+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\left(\Lambda_{i,jk}-\Lambda_{i,(n+1)(n+1)}\delta_{jk}\right)u_{j}x_{k}$
$\equiv$ $(L+1)\Lambda_{i,(n+1)(n+1)}+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}u_{j}x_{k}~{}.$
For simplicity, we denote
$I_{0}(f)\equiv\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}f(x)\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},$
with $f(x)$ a polynomial in $x$. Then we have
$-(L+1)\lambda_{0}\Lambda_{i,(n+1)(n+1)}I_{0}(1)$
$=$ $\lambda_{0}\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}I_{0}(u_{j}x_{k})$
$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\frac{\partial\mathcal{F}^{\lambda_{0}}}{\partial x_{j}}x_{k}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)$
$=$ $-\sum_{j=1}^{n}\Lambda^{\prime}_{i,jj}I_{0}(1)-i\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}I_{0}(x_{k})$
$=$ $(n-n_{0})\Lambda_{i,(n+1)(n+1)}I_{0}(1)-\sum_{j=1}^{n+1}\sum_{k=1}^{n}v_{j}\frac{\partial I_{0}(1)}{\partial v_{k}}~{}.$
That is
$\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}\frac{\partial I_{0}(1)}{\partial v_{k}}=$ $\left[n-n_{0}+(L+1)\lambda_{0}\right]$
×$\times\Lambda_{i,(n+1)(n+1)}I_{0}(1)~{}.$
Thus
$\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}\frac{\partial}{\partial v_{k}}\left[\mathcal{G}(v)^{\lambda_{0}+\frac{n-n_{0}}{L+1}}I_{0}(1)\right]=0~{}.$
By assumption, the matrix $\Lambda^{\prime}_{ij}(v)\equiv\sum_{k=1}^{n+1}\Lambda^{\prime}_{i,kj}v_{k}$ is invertible. Thus $\mathcal{G}(v)^{\lambda_{0}+\frac{n-n_{0}}{L+1}}I_{0}(1)$ is independent of $v_{i},~{}i\neq n+1$. That is
$I_{0}(1)=C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}-\frac{n-n_{0}}{L+1}}~{},$ (A.1)
where $C_{\mathcal{F}}$ is a constant. The $v_{n+1}$ dependence is fixed through a simple power counting.
Appendix B Derivation of eq. (IV.10)
By virtue of eq. IV.5, we have (In this section, summations over repeated indices are always implied.)
$\beta_{ij,k}b_{kl}u_{l}\mathcal{F}=$ $\beta_{ij,k}b_{kl}\frac{\partial\mathcal{F}}{\partial x_{l}}$
$=$ $2\tilde{A}_{ik}B_{kl}p_{l}\cdot p_{j}-\beta_{ij,k}b_{kl}C_{l}U$
$=$ $2\left(A^{-1}\right)_{ik}B_{kl}p_{l}\cdot p_{j}u_{n+1}\mathcal{F}-\beta_{ij,k}b_{kl}C_{l}u_{n+1}\mathcal{F}~{}.$
Thus we have
$B_{ij}=A_{ik}g_{jl}B^{\prime}_{kl}~{}.$
Similarly,
$\alpha_{ij,k}a_{kl}u_{l}\mathcal{F}=\alpha_{ij,k}a_{kl}\frac{\partial\mathcal{F}}{\partial x_{l}}$
$=$ $-U\left[p_{m}\cdot p_{n}\left(A^{-1}\right)_{ik}\left(A^{-1}\right)_{jl}B_{km}B_{ln}+\alpha_{ij,k}a_{kl}C_{l}\right]$
$+\tilde{A}_{ij}\left[p_{k}\cdot p_{l}\left(A^{-1}\right)_{mn}B_{mk}B_{nl}+x_{n+1}-C\right]$
$=$ $-\mathcal{F}u_{n+1}\left(g_{mn}B^{\prime}_{im}B^{\prime}_{jn}+\alpha_{ij,k}a_{kl}C_{l}\right)+\left(A^{-1}\right)_{ij}\mathcal{F}~{}.$
That is
$\left(A^{-1}\right)_{ij}=\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}~{}.$
Taking the determinant of both sides, we get
$\begin{split}\mathcal{G}=&\mathcal{F}^{-1}=u_{n+1}U^{-1}=u_{n+1}\det\left(A^{-1}\right)\\ =&u_{n+1}\det\left[\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{}.\end{split}$ (B.1)
Following some similar algebras, we can also get
$0=$ $A_{ik}\beta_{kj,l}b_{lm}(u_{m}+C_{m}u_{n+1})-2B_{ik}p_{k}\cdot p_{j}u_{n+1}~{},$ (B.2a)
$0=$ $(L+1)\left(A_{ik}\alpha_{kj,l}a_{lm}+\frac{1}{2}B_{ik}\beta_{jk,l}b_{lm}\right)$ (B.2b)
×$\times(u_{m}+C_{m}u_{n+1})-\delta_{ij}u_{k}x_{k}~{}.$ (B.2c)
These equations are of the form of the complete condition eq. III.5, since $A_{ij}$ and $B_{ij}$ are linear in $x$. It is easy to see that they contain $\frac{1}{2}L(L+1)+E=n$ independent equations. $\sum_{j=1}^{n}\Lambda_{i,jj}$ can be obtained from eq. III.5 by replaceing $u_{j}x_{k}$ by $\delta_{jk}$. Thus eq. B.2a gives [Remember the definitions in eq. IV.6.]
$\sum_{j=1}^{n}\Lambda_{i,jj}=\Lambda_{i,(n+1)(n+1)}=0~{}.$
And eq. B.2c gives (Notice that the matrix $A_{ij}$ is of dimension $L\times L$, and the matrix $B_{ij}$ is of dimension $L\times E$.)
$\Lambda_{i,(n+1)(n+1)}=$ $-1~{},$
$\sum_{j=1}^{n}\Lambda_{i,jj}=$ $\frac{1}{2}(L+1)(L+1+E)-n~{}.$
That is,
$n_{0}=n-\frac{1}{2}(L+1)(L+1+E)~{}.$ (B.3)

### v3

@@
 The two representations related through this transformation are dual to each other.
-And it naturally leads to dualities of Landau equations and linear integral relations between the two representations.
+It naturally leads to dualities of Landau equations and linear integral relations between the two representations.
 For integrals with momentum-space correspondences, the dual representation is equivalent to the Baikov representation.
@@
 [24]).
-Furthermore, the parametric representation can reveal some hidden mathematical structures of Feynman integrals due to its connections with hypergeometric functions [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], periods [36, 37, 14], coaction [38, 39, 40, 41, 42], etc.
+Furthermore, the parametric representation can reveal some hidden mathematical structures of Feynman integrals due to its connections with hypergeometric functions [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], periods [36, 37, 14], the coaction [38, 39, 40, 41, 42], etc.
 Another widely used representation is the Baikov representation [43, 44, 45, 46, 47] and its variants [48, 49, 50].
 Although it is less suited for direct integration, the Baikov representation has the advantage that implementing cuts is straightforward [48, 51, 52], and it is more compatible with twisted cohomology theory [53, 54, 55, 56].
-The Baikov representation has been extensively applied to integral reduction [57, 58, 53, 54, 55, 56, 59, 60], constructing differential equation basis [61, 62, 63], Landau analysis [64, 65, 66, 67], etc..
+The Baikov representation has been extensively applied to integral reduction [57, 58, 53, 54, 55, 56, 59, 60], constructing differential equation basis [61, 62, 63], Landau analysis [64, 65, 66, 67], etc.
 While it is evident that these two representations should be equivalent, since both are derived from the momentum representation, it was not clear how to directly derive one representation from the other.
-This problem is addressed in this article.
+This problem is addressed in this letter.
 It is demonstrated that these two representations are related by a simple transformation.
+This letter is organized as follows.
 A transformation on homogeneous polynomials is introduced in section II.
@@
 $\mathcal{G}(u)\equiv\left.\frac{1}{\mathcal{F}}\right|_{x=x(u)}.$ (II.1b)
-We assume that the degree of $\mathcal{F}$ is $L+1$. Then obviously
+We assume that the degree of $\mathcal{F}$ is $L+1$. Then by virtue of the homogeneity of $\mathcal{F}$, we have
 $\sum_{i=1}^{n+1}u_{i}x_{i}=L+1.$ (II.2)
-Taking the derivative of both sides with respect to $u_{i}$, we get
+Taking the derivatives of both sides with respect to $u_{i}$, we get
 $x_{i}+\sum_{j=1}^{n+1}u_{j}\frac{\partial x_{j}}{\partial u_{i}}=0.$
@@
 Even if it is, the obtained function $\mathcal{G}(u)$ may not be a rational function.
-However, for the $\mathcal{F}$ polynomial of a loop integral with a complete set of propagators, $\mathcal{G}$ is indeed a rational function, as will be clear in the subsequent sections.
+However, for the $\mathcal{F}$ polynomial of a loop integral, $\mathcal{G}$ is indeed a rational function, as will be clear in the subsequent sections.
 III Representation transformation
 As was shown in refs. [7, 8, 9], a Feynman integral can be parametrized by scalar integrals of the following form.
-$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})\\ =&\frac{\Gamma(-\lambda_{0})}{\prod_{i=m+1}^{n+1}\Gamma(\lambda_{i}+1)}\int\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\prod_{i=1}^{m}x_{i}^{-\lambda_{i}-1}\prod_{i=m+1}^{n+1}x_{i}^{\lambda_{i}}~{}.\end{split}$ (III.1)
+$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})=\frac{\Gamma(-\lambda_{0})}{\prod_{i=m+1}^{n+1}\Gamma(\lambda_{i}+1)}\\ &\times\int\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\prod_{i=1}^{m}x_{i}^{-\lambda_{i}-1}\prod_{i=m+1}^{n+1}x_{i}^{\lambda_{i}}~{}.\end{split}$ (III.1)
 […]
 We introduce the function [9]
-$w_{\lambda}(x)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}v~{}\frac{1}{(v-i0^{+})^{\lambda+1}}e^{ivx}~{}.$ (III.2)
+$w_{\lambda}(x)\equiv e^{\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}v~{}\frac{1}{(v+i0^{+})^{\lambda+1}}e^{-ivx}~{}.$ (III.2)
 We have
-$w_{-1}(x)=$ $2\pi\delta(x)~{},$
-$w_{\lambda}(x)=$ $2\pi\theta(x)\frac{x^{\lambda}}{\Gamma(\lambda+1)},~{}\lambda\geq 0~{}.$
-And we have
-$\frac{1}{x^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}v}{2\pi}~{}w_{\lambda}(x)e^{ivx}~{}.$ (III.3)
+$w_{-1}(x)=$ $~{}2\pi\delta(x)~{},$
+$w_{\lambda}(x)=$ $~{}2\pi\theta(x)\frac{x^{\lambda}}{\Gamma(\lambda+1)},~{}\lambda\geq 0~{}.$
+The Fourier transform of $w_{\lambda}(x)$ gives
+$\frac{1}{x^{\lambda+1}}=e^{\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}v}{2\pi}~{}w_{\lambda}(v)e^{-ivx},~{}\mathrm{Im}\{x\}<0~{}.$ (III.3)
 Obviously we can write $\frac{1}{\Gamma(\lambda+1)}\int_{0}^{\infty}\mathrm{d}xx^{\lambda}$ as $\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}xw_{\lambda}(x)$.
-By expressing $w_{\lambda}(x)$ (or $x^{-\lambda-1}$) in terms of $v^{-\lambda-1}$ [or $w_{\lambda}(v)$], it is easy to get
-$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}v_{i}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\int_{-\infty}^{\infty}\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\int_{-\infty}^{\infty}\mathrm{d}\Pi_{v}^{(n+1)}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},\end{split}$ (III.4)
-where the last step can be proven by inserting an integral $\int\mathrm{d}s\delta(s-\mathcal{E}^{(1)}(v))$, rescaling $x$ and $v$ by $x_{i}\to\frac{x_{i}}{s}$ and $v_{i}\to sv_{i}$, and finally integrating out $s$.
+By expressing $w_{\lambda}(x)$ (or $x^{-\lambda-1}$) in terms of $v^{-\lambda-1}$ [or $w_{\lambda}(v)$] by using eqs.
+(III.2 and III.3), the integral $I_{x}$ in eq.
+(III.1) becomes
+$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(-2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\\ &\times\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}v_{i}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\\ &\times\int_{-\infty}^{\infty}\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\exp\left(-i\sum_{i=1}^{n+1}v_{i}x_{i}\right)\\ =&(-2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\\ &\times\int_{-\infty}^{\infty}\mathrm{d}\Pi_{v}^{(n+1)}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\\ &\times\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(-i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},\end{split}$ (III.4)
+where $\mathrm{d}\Pi_{v}^{(n+1)}\equiv\prod_{i=1}^{n+1}\mathrm{d}v_{i}~{}\delta(1-\mathcal{E}^{(1)}(v))$.
+The last step of this equation can be proven by inserting an integral $\int\mathrm{d}s\delta(s-\mathcal{E}^{(1)}(v))$, rescaling $x$ and $v$ by $x_{i}\to\frac{x_{i}}{s}$ and $v_{i}\to sv_{i}$, and finally integrating out $s$.
 We consider a special class of integrals of which the $\mathcal{F}$ polynomials are complete. Here, by complete, we mean that there are $n$ linearly independent matrices $\Lambda_{i,jk}$ such that
@@
 We further assume that
-$\Lambda_{i,j(n+1)}=$ $0,~{}j\neq n+1~{},$ (III.6a)
-$\sum_{j=1}^{n}\Lambda_{i,jj}=$ $n_{0}\Lambda_{i,(n+1)(n+1)}~{},$ (III.6b)
+$\Lambda_{i,j(n+1)}=$ $~{}0,~{}j\neq n+1~{},$ (III.6a)
+$\sum_{j=1}^{n}\Lambda_{i,jj}=$ $~{}n_{0}\Lambda_{i,(n+1)(n+1)}~{},$ (III.6b)
 with $n_{0}$ a constant.
-As can be seen in appendix B, normal loop integrals do satisfy these conditions.
-The first condition is a consequence of the fact that $\mathcal{F}$ is linear in $x_{n+1}$.
-For a complete homogeneous $\mathcal{F}$ polynomial satisfying the conditions in eq.
-III.6, it can be proven that (see appendix A)
-$\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)=C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}^{\prime}}~{},$ (III.7)
-where $\lambda_{0}^{\prime}\equiv\lambda_{0}+\frac{n-n_{0}}{L+1}$ and $C_{\mathcal{F}}$ is a constant. Thus we get
-$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)C_{\mathcal{F}}\\ &\times\int_{-\infty}^{\infty}\mathrm{d}\Pi_{u}^{(n+1)}\mathcal{G}(u)^{-\lambda_{0}^{\prime}}\prod_{i=1}^{m}w_{\lambda_{i}^{\prime}}(u_{i})\prod_{i=m+1}^{n+1}u_{i}^{-\lambda_{i}^{\prime}-1}\\ \equiv&(2i\pi)^{-n-1}\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)C_{\mathcal{F}}\frac{\Gamma(-\lambda_{0})}{\Gamma(\lambda_{0}^{\prime})}\\ &\times I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})~{},\end{split}$ (III.8)
-where $\lambda_{n+1}^{\prime}=\lambda_{n+1}+n_{0}+1$, and $\lambda_{i}^{\prime}=\lambda_{i},~{}i\neq 0,n+1$.
+As can be seen in appendix B, a family of loop integrals with a complete set of propagators do satisfy these conditions.
+That is, the completeness of propagators implies the completeness of the corresponding $\mathcal{F}$ polynomial.
+The first condition of eqs.
+(III.6) is a consequence of the fact that $\mathcal{F}$ is linear in $x_{n+1}$.
+For a complete homogeneous $\mathcal{F}$ polynomial satisfying the conditions in eqs.
+(III.6), it can be proven that (see appendix A)
+$\begin{split}&\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)\\ =&~{}(-2i\pi)^{n+1}e^{\frac{i}{2}\pi(L+1)\lambda_{0}}C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{\lambda_{0}^{\prime}}~{},\end{split}$ (III.7)
+where $\lambda_{0}^{\prime}\equiv-\lambda_{0}-\frac{n-n_{0}}{L+1}$ and $C_{\mathcal{F}}$ is a ($v$-independent) constant. Substituting eq. (III.7) into eq. (III.4) gives
+$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&\exp\left(\frac{i\pi}{2}\left[(L+1)\lambda_{0}+\sum_{i=1}^{n+1}\lambda_{i}\right]\right)\\ &\times C_{\mathcal{F}}\frac{\Gamma(-\lambda_{0})}{\Gamma(-\lambda_{0}^{\prime})}~{}I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})~{},\end{split}$ (III.8)
+where $\lambda_{n+1}^{\prime}=\lambda_{n+1}+n_{0}+1$, and $\lambda_{i}^{\prime}=\lambda_{i},~{}i\neq 0,n+1$. The integral $I_{u}$ is defined by
+$I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ (III.9)
+$\equiv$ $~{}\Gamma(-\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\mathrm{d}\Pi_{u}^{(n+1)}\mathcal{G}(u)^{\lambda_{0}^{\prime}}\prod_{i=1}^{m}w_{\lambda_{i}^{\prime}}(u_{i})\prod_{i=m+1}^{n+1}u_{i}^{-\lambda_{i}^{\prime}-1}~{}.$
 Thus, we have proven that $I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})$ equals $I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ up to some constants, and there is an elegant duality between these two representations, which is summarized in table 1.
@@
 $u=\frac{1}{\mathcal{F}}\frac{\partial\mathcal{F}}{\partial x}$ $x=\frac{1}{\mathcal{G}}\frac{\partial\mathcal{G}}{\partial u}$
-$\frac{1}{u^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{2\pi}~{}w_{\lambda}(x)e^{iux}$  ${}w_{\lambda}(x)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}u~{}\frac{1}{u^{\lambda+1}}e^{iux}$
-Table 1: The duality between the two representations.
+$\frac{1}{u^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{2\pi}~{}w_{\lambda}(x)e^{iux}$  ${}w_{\lambda}(x)=e^{\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}u~{}\frac{1}{u^{\lambda+1}}e^{-iux}$
+Table 1: The duality between the two representations. Here the $\frac{1}{u^{\lambda+1}}$ should be understood as $\frac{1}{(u+i0^{+})^{\lambda+1}}$.
 IV The Baikov representation
-We consider a loop integral with a complete set of propagators
+We consider a loop integral with propagators
 $v_{i}=\sum_{j,k}A_{i,jk}l_{j}\cdot l_{k}+2\sum_{j,k}B_{i,jk}l_{j}\cdot p_{k}+C_{i},\quad i=1,~{}2,\dots,n~{},$ (IV.1)
 where $l$ are the loop momenta and $p$ are the external momenta.
-For simplicity, I assume that all the propagators are linearly independent and the Gram determinant of the external momenta $p$ does not vanish.
+For simplicity, we assume that the Gram determinant of the external momenta $p$ does not vanish, and all the propagators are complete and linearly independent.
+Here, by complete, we mean that $v_{i}$ form a complete basis such that all the scalar products $l_{i}\cdot l_{j}$ and $l_{i}\cdot p_{j}$ can be expressed in terms of $v_{i}$.
 For future convenience, we collectively denote $l$ and $p$ by $q$.
 Scalar integrals can be parametrized by integrals of the form [44, 45]
-$J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})=\int\prod_{i=1}^{n}\mathrm{d}v_{n}~{}P^{-\lambda_{0}-\frac{1}{2}(L+E+1)}\prod_{i=1}^{n}v_{i}^{-\lambda_{i}-1}~{},$ (IV.2)
+$\begin{split}&J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})=\\ &\int\prod_{i=1}^{n}\mathrm{d}v_{n}~{}P^{-\lambda_{0}-\frac{1}{2}(L+E+1)}\prod_{i=1}^{n}v_{i}^{-\lambda_{i}-1}~{},\end{split}$ (IV.2)
 where $E$ is the number of the external momenta $p$.
 Here we have omitted some irrelevant prefactors for brevity.
+Notice that we use such a convention that the indices of the propagators are $-\lambda_{i}-1$ rather than the more conventional $-\lambda_{i}$.
 The polynomial $P(v)$ is proportional to the Gram determinant of the momenta $q$.
-That is
+That is,
 $P(v)\equiv\frac{\det(q_{i}\cdot q_{j})}{\det(p_{i}\cdot p_{j})}~{}.$ (IV.3)
 In this section, we will prove that $J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})$ is equivalent to the integral $I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ in eq.
-III.8 up to some constants.
-Taking $\mathcal{E}^{(1)}=|u_{n+1}|$ and $m=0$ in the integral $I_{u}$, we get
+III.9 up to some constants.
+Taking $\mathcal{E}^{(1)}(u)=|u_{n+1}|$ and $m=0$ in the integral $I_{u}$, we get
 $I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$
-$=$ $\Gamma(\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\left[\mathcal{G}_{+}^{-\lambda_{0}^{\prime}}+\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n}^{\prime}+1)}\right]\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
-$=$ $\Gamma(\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
-×$\times\left[\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{-i\pi(\lambda_{n}^{\prime}+1)}+\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n}^{\prime}+1)}\right]$
-$=$ $2\Gamma(\lambda_{0}^{\prime})\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
-$=$ $\frac{-4i}{\Gamma(1-\lambda_{0}^{\prime})}\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)$
-×$\times\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\theta\left(-\mathcal{G}_{-}\right)\left(-\mathcal{G}_{-}\right)^{-\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}~{},$
+$=$ $\Gamma(-\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\left[\mathcal{G}_{+}^{\lambda_{0}^{\prime}}+\mathcal{G}_{-}^{\lambda_{0}^{\prime}}e^{-i\pi(\lambda_{n+1}^{\prime}+1)}\right]\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
+$=$ $\Gamma(-\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
+×$\times\left[\mathcal{G}_{-}^{\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n+1}^{\prime}+1)}+\mathcal{G}_{-}^{\lambda_{0}^{\prime}}e^{-i\pi(\lambda_{n+1}^{\prime}+1)}\right]$
+$=$ $2\Gamma(-\lambda_{0}^{\prime})\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\mathcal{G}_{-}^{\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
+$=$ $\frac{-4i\pi}{\Gamma(1+\lambda_{0}^{\prime})}\cos\left(\pi\lambda_{n+1}^{\prime}\right)$
+×$\times\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\theta\left(-\mathcal{G}_{-}\right)\left(-\mathcal{G}_{-}\right)^{\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}~{},$
 where $\mathcal{G}_{+}\equiv\left.\mathcal{G}\right|_{u_{n+1}=1}$ and $\mathcal{G}_{-}\equiv\left.\mathcal{G}\right|_{u_{n+1}=-1}$.
-Due to the positivity of $|u_{n+1}|$, the integral recieves contributions from both $u_{n+1}=1$ and $u_{n+1}=-1$.
-The second step is carried out by reversing the path of integration for all $x_{i}$ for the first term in the squared bracket.
-And in the last step, we have closed the contour around the branch cut.
+Due to the positivity of $|u_{n+1}|$, the integral receives contributions from both $u_{n+1}=1$ and $u_{n+1}=-1$.
+The second step of this equation is carried out by replacing the integration variables with $x_{i}\to-x_{i}$ for the first term in the square bracket.
+In the last step, we have closed the contour around the branch cut.
+As a result, only the discontinuity of $\mathcal{G}_{-}^{\lambda_{0}^{\prime}}$ contributes to the integral.
+The discontinuity arises from the region where $\mathcal{G}_{-}<0$, which gives rise to the theta function $\theta\left(-\mathcal{G}_{-}\right)$.
 Notice that we use such a convention that $\mathrm{Im}\{\mathcal{F}\}<0$.
-Thus $\mathrm{Im}\{\mathcal{G}\}>0$ when $\mathrm{Re}\{\mathcal{G}\}<0$.
-Now the region of integration is constrained by the theta function $\theta\left(-\mathcal{G}_{-}\right)$.
-Since $n_{0}=n-\frac{1}{2}(L+1)(L+E+1)$ (see appendix B), we have $\lambda_{0}^{\prime}=\lambda_{0}+\frac{1}{2}(L+E+1)$.
+Thus $\mathrm{Im}\{\mathcal{G}\}>0$.
+In the last step we have expressed $\sin(\pi\lambda_{0}^{\prime})$ in terms of $\frac{-\pi}{\Gamma(-\lambda_{0}^{\prime})\Gamma(1+\lambda_{0}^{\prime})}$.
+Since $n_{0}=n-\frac{1}{2}(L+1)(L+E+1)$ (see appendix B), we have $\lambda_{0}^{\prime}=-\lambda_{0}-\frac{1}{2}(L+E+1)$.
 Comparing the above integral with the Baikov integral in eq.
 IV.2, we see that these two integrals are equivalent if
-$P(v_{1},v_{2},\dots,v_{n})=-\mathcal{G}_{-}=-\mathcal{G}(v_{1},v_{2},\dots,v_{n},-1)~{}.$ (IV.4)
+$\begin{split}P(v_{1},v_{2},\dots,v_{n})=&-\mathcal{G}_{-}\\ =&-\mathcal{G}(v_{1},v_{2},\dots,v_{n},-1)~{}.\end{split}$ (IV.4)
 Immediately, we will see that this is indeed true.
@@
 Let $a_{ij}$, $b_{ij}$, $\alpha_{ij,k}$, and $\beta_{ij}$ be those defined in ref. [8]. That is, they are solutions to the following equations.
-$\sum_{l}a_{il}B_{l,jk}=0,$ (IV.6)
-$\sum_{l}b_{il}A_{l,jk}=0,$ (IV.7)
-$\sum_{kl}\alpha_{ij,k}a_{kl}A_{l,mn}=\frac{1}{2}\left(\delta^{im}\delta^{jn}+\delta^{in}\delta^{jm}\right),$ (IV.8)
-$\sum_{kl}\beta_{ij,k}b_{kl}B_{l,mn}=\delta^{im}\delta^{jn}~{}.$ (IV.9)
+$\sum_{l}a_{il}B_{l,jk}=0,$ (IV.6a)
+$\sum_{l}b_{il}A_{l,jk}=0,$ (IV.6b)
+$\sum_{kl}\alpha_{ij,k}a_{kl}A_{l,mn}=\frac{1}{2}\left(\delta^{im}\delta^{jn}+\delta^{in}\delta^{jm}\right),$ (IV.6c)
+$\sum_{kl}\beta_{ij,k}b_{kl}B_{l,mn}=\delta^{im}\delta^{jn}~{}.$ (IV.6d)
 Then it can be proven that (see appendix B)
-$\begin{split}\mathcal{G}=&u_{n+1}\det\left[\sum_{kl}\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\right.\\ &\left.+u_{n+1}\sum_{mn}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{},\end{split}$ (IV.10)
+$\begin{split}\mathcal{G}=~{}&u_{n+1}\det\left[\sum_{kl}\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\right.\\ &\left.+u_{n+1}\sum_{mn}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{},\end{split}$ (IV.7)
 where $u_{i}$ are those appeared in eq. II.1, $g_{ij}$ is the inverse of the Gram matrix $p_{i}\cdot p_{j}$, and
@@
 On the other hand, $q_{i}\cdot q_{j}$ can be understood as a block matrix. Then we have
-$P=$ $\frac{1}{\det(p_{i}\cdot p_{j})}\det\begin{pmatrix}l_{i}\cdot l_{j}&l_{i}\cdot p_{j}\\ l_{i}\cdot p_{j}&p_{i}\cdot p_{j}\end{pmatrix}$
-$=$ $\det\left(l_{i}\cdot l_{j}-\sum_{mn}g_{mn}l_{i}\cdot p_{m}l_{j}\cdot p_{n}\right)~{}.$
-It is easy to get
+$P=~{}$ $\frac{1}{\det(p_{i}\cdot p_{j})}\det\begin{pmatrix}l_{i}\cdot l_{j}&l_{i}\cdot p_{j}\\ l_{i}\cdot p_{j}&p_{i}\cdot p_{j}\end{pmatrix}$
+[…]
+Solving eq. (IV.1) for $l_{i}\cdot l_{j}$ and $l_{i}\cdot p_{j}$, we get
 $l_{i}\cdot p_{j}=$ $\frac{1}{2}\sum_{k,l}\beta_{ij,k}b_{kl}(v_{l}-C_{l})\equiv-B^{\prime\prime}_{ij}=-\left.B^{\prime}_{ij}(v)\right|_{v_{n+1}=-1}~{},$
@@
 Thus
-$P=\det\left(\sum_{kl}\alpha_{ij,k}a_{kl}(v_{l}-C_{l})-\sum_{mn}g_{mn}B^{\prime\prime}_{im}B^{\prime\prime}_{jn}\right)~{}.$ (IV.11)
-Comparing eq. IV.10 with eq. IV.11, we get eq. IV.4.
+$\begin{split}P=&\det\left(\sum_{kl}\alpha_{ij,k}a_{kl}(v_{l}-C_{l})\right.\\ &\left.-\sum_{mn}g_{mn}B^{\prime\prime}_{im}B^{\prime\prime}_{jn}\right)~{}.\end{split}$ (IV.8)
+Comparing eq. IV.7 with eq. IV.8, we get eq. IV.4.
 V Discussion
@@
 Furthermore, it is proven that the dual representation is equivalent to the Baikov representation.
-The transformation and the corresponding duality can be applied to various aspects of Feynman integrals, including Laundau analysis and integral reduction.
-Specifically, the transformation in II.1 establishes a duality between the first (second) Landau equation in the parametric representation and the second (first) Landau equation in the Baikov representation.
-Additionally, eq.
+The transformation and the corresponding duality can be widely applied to the studies of Feynman integrals, including integral reduction and Laundau analysis.
+Specifically, eq.
 III.5 establishes a correspondence between a syzygy equation in the Baikov representation [57, 58] and a parametric annihilator in the parametric representation [6, 17].
+Additionally, the transformation in II.1 establishes a duality between the first (second) Landau equation in the parametric representation and the second (first) Landau equation in the Baikov representation.
+It is known that the infrared singularities of Feynman integrals can be identified with solutions of Landau equations.
+The method of regions [68, 69, 70, 71, 72] provides a systematic method to isolate the infrared structures of Feynman integrals.
+While endpoint singularities (in the parametric space) can be identified using a geometric method [71, 73, 74], this method can not directly be applied to more general regions, including threshold regions [68, 75, 76, 77].
+The latter are related to solutions of the second Landau equations which are away from the origin.
+By virtue of the duality found in this letter, it is expected that this kind of regions can be identified by applying the geometric method to the dual representation.
+This is confirmed through some one-loop integrals, including those considered in refs.
+[68, 75].
+The application of this method to multiloop integrals is complicated by the presence of irreducible scalar products, which deserves further investigation.
 Acknowledgements.
@@
 Combining the homogeneity condition eq. II.2 with those conditions in eqs. III.5 and III.6, we have
-$0=$ $\sum_{j,k=1}^{n+1}\Lambda_{i,jk}u_{j}x_{k}$
-$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda_{i,jk}u_{j}x_{k}+\Lambda_{i,(n+1)(n+1)}u_{n+1}x_{n+1}$
-$=$ $(L+1)\Lambda_{i,(n+1)(n+1)}$
-$+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\left(\Lambda_{i,jk}-\Lambda_{i,(n+1)(n+1)}\delta_{jk}\right)u_{j}x_{k}$
-$\equiv$ $(L+1)\Lambda_{i,(n+1)(n+1)}+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}u_{j}x_{k}~{}.$
+$\begin{split}0=&\sum_{j,k=1}^{n+1}\Lambda_{i,jk}u_{j}x_{k}\\ =&\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda_{i,jk}u_{j}x_{k}+\Lambda_{i,(n+1)(n+1)}u_{n+1}x_{n+1}\\ =&(L+1)\Lambda_{i,(n+1)(n+1)}+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}u_{j}x_{k}~{},\end{split}$ (A.1)
+with
+$\Lambda^{\prime}_{i,jk}\equiv\Lambda_{i,jk}-\Lambda_{i,(n+1)(n+1)}\delta_{jk}~{}.$
 For simplicity, we denote
-$I_{0}(f)\equiv\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}f(x)\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},$
-with $f(x)$ a polynomial in $x$. Then we have
+$I_{0}(f)\equiv\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}f(x)\exp\left(-i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},$
+with $f(x)$ a polynomial in $x$. Then, by virtue of eq. (A.1), we have
 $-(L+1)\lambda_{0}\Lambda_{i,(n+1)(n+1)}I_{0}(1)$
 $=$ $\lambda_{0}\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}I_{0}(u_{j}x_{k})$
-$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\frac{\partial\mathcal{F}^{\lambda_{0}}}{\partial x_{j}}x_{k}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)$
-$=$ $-\sum_{j=1}^{n}\Lambda^{\prime}_{i,jj}I_{0}(1)-i\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}I_{0}(x_{k})$
+$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\frac{\partial\mathcal{F}^{\lambda_{0}}}{\partial x_{j}}x_{k}\exp\left(-i\sum_{i=1}^{n+1}v_{i}x_{i}\right)$
+$=$ $-\sum_{j=1}^{n}\Lambda^{\prime}_{i,jj}I_{0}(1)+i\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}I_{0}(x_{k})$
 $=$ $(n-n_{0})\Lambda_{i,(n+1)(n+1)}I_{0}(1)-\sum_{j=1}^{n+1}\sum_{k=1}^{n}v_{j}\frac{\partial I_{0}(1)}{\partial v_{k}}~{}.$
@@
 ×$\times\Lambda_{i,(n+1)(n+1)}I_{0}(1)~{}.$
-Thus
+On the other hand, eq. (A.1) is equivalent to
+$\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}\frac{\partial\mathcal{G}(v)}{\partial v_{k}}=-(L+1)\Lambda_{i,(n+1)(n+1)}\mathcal{G}~{}.$
+Thus we get
 $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}\frac{\partial}{\partial v_{k}}\left[\mathcal{G}(v)^{\lambda_{0}+\frac{n-n_{0}}{L+1}}I_{0}(1)\right]=0~{}.$
@@
 That is
-$I_{0}(1)=C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}-\frac{n-n_{0}}{L+1}}~{},$ (A.1)
-where $C_{\mathcal{F}}$ is a constant. The $v_{n+1}$ dependence is fixed through a simple power counting.
-Appendix B Derivation of eq. (IV.10)
+$\begin{split}I_{0}(1)=&(-2i\pi)^{n+1}e^{\frac{i}{2}\pi(L+1)\lambda_{0}}C_{\mathcal{F}}\\ &\times v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}-\frac{n-n_{0}}{L+1}}~{},\end{split}$ (A.2)
+where $C_{\mathcal{F}}$ is a constant, and the factors in $i$ and $\pi$ are added for future convenience. The $v_{n+1}$ dependence is fixed through a simple power counting.
+Appendix B Derivation of eq. (IV.7)
 By virtue of eq. IV.5, we have (in this section, summations over repeated indices are always implied.)
@@
 $=$ $2\tilde{A}_{ik}B_{kl}p_{l}\cdot p_{j}-\beta_{ij,k}b_{kl}C_{l}U$
-$=$ $2\left(A^{-1}\right)_{ik}B_{kl}p_{l}\cdot p_{j}u_{n+1}\mathcal{F}-\beta_{ij,k}b_{kl}C_{l}u_{n+1}\mathcal{F}~{}.$
-Thus we have
-$B_{ij}=A_{ik}g_{jl}B^{\prime}_{kl}~{}.$
+$=$ $2\left(A^{-1}\right)_{ik}B_{kl}p_{l}\cdot p_{j}u_{n+1}\mathcal{F}$
+$-\beta_{ij,k}b_{kl}C_{l}u_{n+1}\mathcal{F}~{}.$
+Multiplying both sides with $\frac{1}{2u_{n+1}\mathcal{F}}A_{ij}g_{kl}$, we get
+$B_{ij}=A_{ik}g_{jl}B^{\prime}_{kl}~{}.$ (B.1)
 Similarly,
@@
 That is
-$\left(A^{-1}\right)_{ij}=\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}~{}.$
+$\begin{split}\left(A^{-1}\right)_{ij}=&\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\\ &+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}~{}.\end{split}$ (B.2)
 Taking the determinant of both sides, we get
-$\begin{split}\mathcal{G}=&\mathcal{F}^{-1}=u_{n+1}U^{-1}=u_{n+1}\det\left(A^{-1}\right)\\ =&u_{n+1}\det\left[\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{}.\end{split}$ (B.1)
+$\begin{split}\mathcal{G}=&\mathcal{F}^{-1}=u_{n+1}U^{-1}=u_{n+1}\det\left(A^{-1}\right)\\ =&u_{n+1}\det\left[\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\right.\\ &\left.+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{}.\end{split}$ (B.3)
 Following some similar algebras, we can also get
-$0=$ $A_{ik}\beta_{kj,l}b_{lm}(u_{m}+C_{m}u_{n+1})-2B_{ik}p_{k}\cdot p_{j}u_{n+1}~{},$ (B.2a)
-$0=$ $(L+1)\left(A_{ik}\alpha_{kj,l}a_{lm}+\frac{1}{2}B_{ik}\beta_{jk,l}b_{lm}\right)$ (B.2b)
-×$\times(u_{m}+C_{m}u_{n+1})-\delta_{ij}u_{k}x_{k}~{}.$ (B.2c)
+$0=$ $A_{ik}\beta_{kj,l}b_{lm}(u_{m}+C_{m}u_{n+1})$
+$-2B_{ik}p_{k}\cdot p_{j}u_{n+1}~{},$ (B.4a)
+$0=$ $(L+1)\left(A_{ik}\alpha_{kj,l}a_{lm}+\frac{1}{2}B_{ik}\beta_{jk,l}b_{lm}\right)$
+×$\times(u_{m}+C_{m}u_{n+1})-\delta_{ij}u_{k}x_{k}~{}.$ (B.4b)
+Here, the first equation is obtained from eq.
+(B.1) by multiplying both sides with $u_{n+1}$, and the second equation is obtained from eq.
+(B.2) by multiplying both sides with the matrix $A_{ij}$ and replacing $\delta_{ij}$ by $\frac{1}{L+1}\delta_{ij}u_{k}x_{k}$.
 These equations are of the form of the complete condition eq.
 III.5, since $A_{ij}$ and $B_{ij}$ are linear in $x$.
-It is easy to see that they contain $\frac{1}{2}L(L+1)+E=n$ independent equations.
+It is easy to see that they contain $\frac{1}{2}L(L+1)+LE=n$ independent equations.
 $\sum_{j=1}^{n}\Lambda_{i,jj}$ can be obtained from eq.
-III.5 by replaceing $u_{j}x_{k}$ by $\delta_{jk}$.
+III.5 by replacing $u_{j}x_{k}$ by $\delta_{jk}$.
 Thus eq.
-B.2a gives [Remember the definitions in eq.
-IV.6.]
-$\sum_{j=1}^{n}\Lambda_{i,jj}=\Lambda_{i,(n+1)(n+1)}=0~{}.$
-And eq. B.2c gives (Notice that the matrix $A_{ij}$ is of dimension $L\times L$, and the matrix $B_{ij}$ is of dimension $L\times E$.)
+B.4a gives [remember the definitions in eqs.
+(IV.6).]
+$\sum_{j=1}^{n}\Lambda_{i,jj}=\Lambda_{i,(n+1)(n+1)}=0~{},$
+and eq. B.4b gives (notice that the matrix $A_{ij}$ is of dimension $L\times L$, and the matrix $B_{ij}$ is of dimension $L\times E$.)
 $\Lambda_{i,(n+1)(n+1)}=$ $-1~{},$